if it takes sally 4 hours to paint a house and it take john 6 hours to paint the same house, then how long will it take both of them at the same time?
Answer by Paul(988)

Sally paints 1/4 of house in one hour
John paints 1/6 of house in one hour
1/4 + 1/6 = 3/12 + 2/12

5/12 = 12/5 = 2.4
Hence together they paint the house in 2.4 hours.